tutor I am having difficulty with this math problem
Solve (2x-5)(x+8)=0

Answer by oscargut(2103)

(2x-5)(x+8)=0
then 2x-5=0 or x+8=0
then x=5/2 or x=-8
Answer: solutions are x=5/2 and x=-8